[Brahms-l] Shift report 20050426 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Apr 26 2005 - 07:58:40 EDT
Shifter leader was RP
Shifters were 


one store till 3:00
pedestal and TPC pedestal runs
beam on again at 5:00
FS at 8 deg , 2442A,
MRS at 45 deg, 1050B

statistic for trigger 2: 3420.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Apr 26 07:59:18 2005

This archive was generated by hypermail 2.1.8 : Tue Apr 26 2005 - 07:59:32 EDT